OpenStudy (anonymous):

What is the measure of an angle whose sine is square root of 3 over 2? 60 45 30

OpenStudy (anonymous):

well it's 60

OpenStudy (anonymous):

sin(x) = sqrt(3)/2 sin^-1( (sqrt(3)/2) = x The angle should be on your unit circle

OpenStudy (anonymous):

sin 60 = sqr3 / 2